OpenStudy (gigi1103):

Which ordered pairs are solutions to the equation? y = x – 4 A. (10, 6) and (1, 5) B. (10, 4) and (5, 1) C. (10, 6) and (5, 1) D. (10, 5) and (5, 1)

OpenStudy (gigi1103):

@StudyGurl14

OpenStudy (anonymous):

I would chose C, probably.

OpenStudy (gigi1103):

can you help

OpenStudy (gigi1103):

@ganeshie8

OpenStudy (ashes.boom):

(x, y) is an ordered pair so you can input the choises into the equation to see which one works if you want.

OpenStudy (gigi1103):

i know that but i dont get how with coordinates

OpenStudy (ashes.boom):

You can also graph the line do you know how to do that? The equation is in slope intercept form

OpenStudy (gigi1103):

can you walk me thru it

OpenStudy (ashes.boom):

Of course!

OpenStudy (ashes.boom):

y = x - 4 is the equation in slope intercept form (y=mx + b) the b represents the y intercept which is where the line intersects with the y axis

OpenStudy (gigi1103):

umm i get the x and y axis part

OpenStudy (gigi1103):

but how do i find it using subtraction with coordinates

OpenStudy (ashes.boom):

Just a second ill be right back

OpenStudy (gigi1103):

ok

OpenStudy (gigi1103):

@BlossomCake

OpenStudy (anonymous):

When you think about it, it's pretty simple. Just plug in the coordinates to the given equation :)

OpenStudy (gigi1103):

but the part i dont get is how to subtract 4 from coordinates

OpenStudy (gigi1103):

ya know?^_^

OpenStudy (gigi1103):

@bohotness

OpenStudy (gigi1103):

@NeonStrawsForever

OpenStudy (anonymous):

For example, in A. the given coordinates are (10, 6) and (1, 5), so to plug them in you would get: 6 = 10 - 4 How? Because your given equation is y = x - 4, where 10 is x and 6 is y. So, then you can ask yourself is six really equal to 10 - 4, the answer is yes. So then you can plug in (1, 5) to get: 5 = 1 - 4 Again, ask yourself, is 1 - 4 really equal to 5? The answer is no, because 1 - 4 equals -3. So, now we know that option A. is NOT an answer, because both coordinates are not true for the equation, only one of them is.

OpenStudy (anonymous):

Now, can you try to plug in the coordinates for B. in (individually) and tell me if they are correct or not?

OpenStudy (gigi1103):

omg that makes sense one sec

OpenStudy (anonymous):

And sorry for the wait!

OpenStudy (anonymous):

So, is B. a possible answer? @gigi1103

OpenStudy (gigi1103):

yes!! am i right

OpenStudy (anonymous):

Nope, because, although (5, 1) proves to be true, (10, 4) doesn't. That means it has to be C. or D. Can you figure them out and tell me which one it is?

OpenStudy (anonymous):

Also, if you don't understand let me know and we can go over it again :)

OpenStudy (gigi1103):

ok!!! lemme think

OpenStudy (anonymous):

And remember that both coordinates must prove to be true in the equations, not just one.

OpenStudy (gigi1103):

d!!

OpenStudy (anonymous):

Okay, so what did you do to get the answer?

OpenStudy (gigi1103):

i plugged in both

OpenStudy (gigi1103):

:)

OpenStudy (anonymous):

And were they both true?

OpenStudy (anonymous):

Remember that our equation is y = x - 4 and the coordinates are (10, 5) and (5, 1).

OpenStudy (gigi1103):

yup :)

OpenStudy (anonymous):

So, for the first one it would be: 5 = 10 - 4 , is this equation true or false?

OpenStudy (anonymous):

How about this one? 1 = 5 - 4

OpenStudy (anonymous):

No, the answer is not true, does 10 - 4 = 5 ? No, 10 - 4 = 6

OpenStudy (gigi1103):

the first ones false

OpenStudy (anonymous):

And for 1 = 5 - 4 , well, this one is true because 5 - 4 is 1.

OpenStudy (anonymous):

Yep. How about 1 = 5 - 4 ?

OpenStudy (gigi1103):

and the second ones truw

OpenStudy (anonymous):

Good, so does this make D. the right answer or not?

OpenStudy (gigi1103):

thats true

OpenStudy (gigi1103):

soo its c

OpenStudy (anonymous):

@gigi1103